how many liters of a 30% solution should be added to 50 liters of a 15% solution if the final mixture should contain 20% solution?
Answer by stanbon(75887)

how many liters of a 30% solution should be added to 50 liters of a 15% solution if the final mixture should contain 20% solution?
------------------------
Equation:
active + active = active
0.30x + 0.15*50 = 0.20(x+50)
---
Multiply thru by 100 to get:
30x + 15*50 = 20x + 20*50
10x = 5*50
x = 25 liters (amt. of 30% solution needed in the mixture)
===============
Cheers,
Stan H.